Problem A.1


The graph below is made of three line segments:


-1 1 2 3 4 5 6 7 8 9 10 11 12


1


2


3


4


y


x


f(x)


g(x)


h(x)


The segments correspond to the following three functions:


f(x) = x − 2, g(x) = p


4 − (x − 6)2 + 2, h(x) = x − 6


Find the total length L of the graph between x = 2 and x = 10.


1
Expert's answer
2021-11-23T11:59:27-0500

"f(x) = x \u2212 2" for "x\isin [2,4]"

"g(x) = \sqrt{4 − (x − 6)^2} + 2" for "x\isin [4,8]"

"h(x) = x \u2212 6" for "x\isin [8,10]"


then:

length of f(x):

"L_1=\sqrt{2^2+2^2}=2\sqrt 2"


length of h(x):

"L_2=\sqrt{2^2+2^2}=2\sqrt 2"


length of g(x):

"L_3=2\pi R/2=2\pi"


total length:

"L=L_1+L_2+L_3=2\sqrt 2+2\sqrt 2+2\pi=2(\pi+2\sqrt 2)"